Tyler Tooling Company uses a job order cost system with overhead applied to products on the basis of machine hours. For the upcoming year, the company estimated its total manufacturing overhead cost at $249,600 and total machine hours at 62,400. During the first month of operations, the company worked on three jobs and recorded the following actual direct materials cost, direct labor cost, and machine hours for each jot Job 101 Job 102 S 8,400 S 5,000 Job 103 S 5,800 S 4,800 Total Direct materials used Direct labor Machine hours $ 10,900 $17,000 S 25,100 26,800 1,000 hours 2,900 hours 900 hours 4,800 hours Job 101 was completed and sold for $50,600 Job 102 was completed but not sold. Job 103 is still in process Actual overhead costs recorded during the first month of operations totaled $14,200 Required 1. Prepare a journal entry showing the transfer of Job 102 into Finished Goods Inventory upon its completion. 2. Prepare the journal entries to recognize the sales revenue and cost of goods sold for Job 101 3. Prepare the journal entry to transfer the balance of the Manufacturing Overhead account to Cost of Goods Sold (If no entry is required for a transaction/event, select No Journal Entry Required in the first account field. Do not round your intermediate calculations.)
Answer #1

Overhead rate = 249600/62400 = $4 per machine hour

a) Journal entries

Date account and explanation debit credit
Finished goods (8400+5000+2900*4) 25000
Work in process 25000
(To record completion of job 102)

b) Journal entries

Date account and explanation debit credit
Cash/Account receivable 50600
Sales revenue 50600
(To record sales)
Cost of goods sold (10900+17000+4000) 31900
Finished goods 31900
(To record cost of good sold)

c) Applied overhead = 4800*4 = 19200; actual overhead = 14200

Over applied overhead = 19200-14200 = $5000

Date account and explanation debit credit
Manufacturing overhead 5000
Cost of goods sold 5000
(To record over applied)